OpenStudy (anonymous):

Find the directional derivaitve of the hemisphere z=sqrt(a^2-x^2-y^2) at the point (a/sqrt(3),a/sqrt(3),a/sqrt(3)) in the direction making an angle pi/4 with the positive x-axis

OpenStudy (lgbasallote):

i am confident that @Kreshnik knows this so i leave you in his most capable hands :P

OpenStudy (anonymous):

lol @lgbasallote you got to be kidding.. :A

OpenStudy (kinggeorge):

There is so much going on in this problem. It scares me.

OpenStudy (anonymous):

well we have a unit vector rightL sqrt(2)/2(i)+sqrt(2)/2(j)

OpenStudy (anonymous):

but this hemisphere is what is bothering me

OpenStudy (anonymous):

how do i compute the gradient

OpenStudy (anonymous):

The directional derivative at a point is gradient(z)(point) dot product with a unit vector in the asking direction.

OpenStudy (anonymous):

\[ \sqrt(2)/2(i)+\sqrt(2)/2(j)\]

OpenStudy (anonymous):

okay, so then that a^2 is a constant number, and we would get a gradeient in the i and j directions only

OpenStudy (anonymous):

They should specify that the vector is in the xy-plane.

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

what do i plug in for a when i get the gradient

OpenStudy (anonymous):

boy, and i thought this place was full of smart people

OpenStudy (anonymous):

so the directional deriavitve is -sqrt(2)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That is the gradient of z \[ \left\{-\frac{x}{\sqrt{a^2-x^2 -y^2}},-\frac{y}{\sqrt{a^2- x^2-y^2}}\right\} \] at the given point \[ \left\{-\frac{a}{\sqrt{a^2}},- \frac{a}{\sqrt{a^2}}\right\} \] Which is {-1,-1}